### v2

# Fuzzy Ontology Embeddings and Visual Query Building for Ontology Exploration

Abstract
Ontologies play a central role in structuring knowledge across domains, supporting tasks such as reasoning, data integration, and semantic search. However, their large size and complexity—particularly in fields such as biomedicine, computational biology, law, and engineering—make them difficult for non-experts to navigate. Formal query languages such as SPARQL offer expressive access but require users to understand the ontology’s structure and syntax. In contrast, visual exploration tools and basic keyword-based search interfaces are easier to use but often lack flexibility and expressiveness.
We introduce FuzzyVis, a proof-of-concept system that enables intuitive and expressive exploration of complex ontologies. FuzzyVis integrates two key components: a fuzzy logic-based querying model built on fuzzy ontology embeddings, and an interactive visual interface for building and interpreting queries. Users can construct new composite concepts by selecting and combining existing ontology concepts using logical operators such as conjunction, disjunction, and negation. These composite concepts are matched against the ontology using fuzzy membership-based embeddings, which capture degrees of membership and support approximate, concept-level similarity search.
The visual interface supports browsing, query composition, and partial search without requiring formal syntax. By combining fuzzy semantics with embedding-based reasoning, FuzzyVis enables flexible interpretation, efficient computation, and exploratory learning. Case studies demonstrate how FuzzyVis supports subtle information needs and helps users uncover relevant concepts in large, complex ontologies.
Keywords: Ontology, Query Building, Ontology Embedding, Visualization, Fuzzy Logic
1 Introduction
Ontologies are formal representations of knowledge that describe the concepts in a knowledge domain, their relationships, and associated instances. They provide a structured framework for representing and reasoning about knowledge, supporting areas such as knowledge representation, artificial intelligence (AI), and data management. In knowledge representation and AI, ontologies help standardize terminology and organize domain knowledge in large datasets, enabling automated reasoning tasks such as inference, problem-solving, and theorem proving [3, 24, 33]. They often serve as advanced thesauri for query expansion, improving search effectiveness and interoperability [6, 4]. In data management, ontologies are central to data integration by supporting the unification of heterogeneous data sources [20] and enhancing concept recognition. A key approach is ontology-based data access, which uses ontologies to enable reasoning over large datasets—typically stored in databases—to answer complex queries [15].
Ontologies are widely used to represent structured knowledge across various domains. In biology, the Gene Ontology (GO) [5] and the Sequence Ontology (SO) [23] provide standardized annotations for genes and sequences. In medicine, the Human Phenotype Ontology (HPO) [29], the Disease Ontology (DO) [52], and MeSH (Medical Subject Headings) [43] support standardized indexing and search of biomedical literature and clinical data. In the legal domain, ontologies such as LKIF-Core [14], LODE [11], and ECLAP [50] support legal reasoning and document management. When no ontology exists for a given domain, one can often be developed to capture its core concepts and relationships.
A persistent challenge in working with ontologies is their size and complexity. Many widely used ontologies—such as GO and HPO—contain hundreds of thousands of concepts, making them difficult to navigate, search, and understand, particularly for non-expert users. Using such ontologies effectively requires exploratory learning, a process through which users build mental models by identifying key concepts, understanding their relationships, and uncovering structural patterns [21]. This process often involves locating landmarks (central or well-known concepts), following routes (chains of related concepts), and exploring neighborhoods (clusters of semantically related concepts).
Exploratory learning over ontologies often involves two core functionalities: searching for relevant concepts and querying to express complex information needs. Search is typically based on keyword matching against concept labels or textual descriptions in the ontology. While simple to use, such methods frequently fail due to vocabulary mismatch, ambiguity, or variation in phrasing. On the other end of the spectrum, formal query languages such as SPARQL [53], OWL-QL [26], SeRQL [13], and SPARQL-DL [54] offer expressive power but require users to understand both the ontology’s structure and the syntax of the query language. These limitations make them difficult to use, especially for non-experts, and the results often require further interpretation. As a result, users exploring large ontologies are left with few practical options when their information needs are vague, evolving, or difficult to articulate precisely.
To address this gap, we propose a querying approach that strikes a balance between expressiveness and simplicity. Instead of writing formal queries, users define new concepts by composing existing primitive concepts in the ontology using set operators—conjunction, disjunction, and negation. These user-defined composite concepts are treated as fuzzy queries and are compared against the ontology to retrieve the most semantically similar primitive concepts. This supports intuitive interaction, flexible query construction, and approximate matching—even when the target composite concept is not explicitly defined in the ontology or when the query is underspecified or partially conflicting. Our approach enables users, particularly non-experts, to express their intent using simple, interpretable building blocks while still benefiting from the structure and semantics of the ontology. We now illustrate this with an example.
Example 1.
To illustrate our querying approach, consider a physician, whom we refer to as the user, exploring HPO to investigate a patient’s symptoms: difficulty speaking and swallowing, with no indication of immune dysfunction. The user suspects a neuromuscular issue but is unsure of the precise terminology.
Through browsing and personal knowledge, the user finds three relevant primitive concepts: Slurred speech (HP_0001350), Dysphagia (difficulty or discomfort in swallowing, HP_0002015), and Abnormality of the immune system (HP_0002715). Using our interface, they define a composite query concept :
This concept captures the co-occurrence of speech and swallowing difficulties while explicitly excluding immune-related causes. Such a combination is not explicitly defined in the ontology and would typically yield no results using standard (non-fuzzy) query tools.
Using our fuzzy querying approach, the system interprets this as a soft concept and returns the most similar known phenotypes. These may include Pseudobulbar paralysis (HP_0007024), referring to a neurological condition, which exhibits some of the patient’s symptoms, or Abnormal esophagus physiology (HP_0025270), referring to structural abnormalities of the patient’s esophagus, which could lead to food entering the air way. Both are relevant concepts to the user’s intent and could be difficult to find through keyword search alone.
To implement our querying approach, we represent ontologies as fuzzy ontologies with fuzzy interpretations, where each concept is treated as a fuzzy set assigning a membership degree to each element in the interpretation domain. Given an ontology, we use existing fuzzy reasoners—such as fuzzyDL [9], DeLorean [48], and FALCON [58]—to construct such interpretations that approximately satisfy the ontology’s axioms. Based on these interpretations, we introduce a novel method for generating ontology embeddings, where each concept is represented as a vector of membership degrees over a fixed, ordered set of domain elements. This allows us to precompute embeddings for all primitive concepts offline. At query time, when a user defines a composite concept using conjunction, disjunction, or negation over known primitives, its embedding is computed by applying the corresponding fuzzy operations element-wise to the membership vectors. This compositional capability enables flexible query construction and is not supported by existing ontology embedding methods. The resulting embeddings also support efficient similarity computations (e.g., cosine similarity) between user-defined concepts and existing primitive concepts in the ontology, enabling approximate matching even when no exact concept exists.
We implement this querying approach in FuzzyVis, a visual system for ontology exploration. In addition to standard features such as keyword search and a typical ontology views, FuzzyVis provides a complex interactive visualization with features such as fisheye distortion and annotation for navigating large ontologies. Users can build visual queries by collecting relevant primitive concepts and combining them using logical operations such as conjunction, disjunction, and negation. The system computes the embedding of the resulting composite concept and returns the most similar primitive concepts as suggestions. Beyond querying, FuzzyVis supports interactive exploration by linking search, query results, and navigation. Concepts returned in results can be located in the primary visualization, and collected concepts can be reused across queries. By combining fuzzy ontology embeddings, query answering, and an intuitive visual interface, FuzzyVis enables non-experts to query and explore large ontologies effectively. It supports exploratory learning through flexible query construction, approximate matching, and structural similarity, addressing key limitations of existing ontology querying tools.
The paper is organized as follows. Section 2 reviews related work. Section 3 introduces the necessary preliminaries and background concepts. Section 4 presents our fuzzy ontology embeddings and describes our querying approach based on these embeddings. Section 5 details the FuzzyVis system, its implementation, and practical use cases. Section 6 concludes the paper and outlines future directions.
2 Related Work
This section reviews related work on ontology embedding methods (Section 2.1) and existing tools for ontology visualization and exploration (Section 2.2).
2.1 Ontology Embeddings
To support scalable computation over ontologies, numerous methods embed ontology elements into continuous spaces that preserve semantic relationships and enable downstream tasks such as classification, clustering, reasoning, and search [18]. Ontology embedding refers to the process of mapping elements of an ontology to structured mathematical representations—typically vectors, regions, or lattices—that reflect their semantics. Formally, an embedding function maps each ontology element (e.g., a concept, relationship, or individual) to a vector in a -dimensional space. Below, we review the main families of embedding methods, organized by their modeling approach:
Translation-based embeddings. These methods model relations as translation operations in vector space. Given a triple —where and are individuals (head and tail), and is a relation—the goal is to learn embeddings such that . This approach was introduced by TransE [12], which assumes relational meaning can be captured by a fixed vector offset. Extensions such as TransH [61] and TransR [41] introduce projection mechanisms to better handle complex relations, embedding entities and relations in different spaces. While scalable and widely used for knowledge graph completion, these models are limited to triple-based structures and struggle with more complex ontology constructs such as concept hierarchies and logical combinations.
Path-based and lexical methods. These methods extract sequential patterns from ontologies or knowledge graphs and apply language modeling techniques to learn embeddings. The key idea is that concepts or entities frequently co-occurring along paths in the graph are likely to be semantically related. RDF2Vec [51] performs random walks over RDF graphs to generate sequences, which are embedded using word2vec. OWL2Vec [16] extends this to OWL ontologies by incorporating logical axioms, syntactic features, and lexical information across multiple ontology projections (e.g., logical, lexical, assertional). These methods are interpretable and scalable, especially for lightweight ontologies with rich textual annotations. However, they often lack alignment with formal semantics and cannot guarantee logical consistency.
Geometric and neurosymbolic embeddings. These methods represent ontology elements as geometric objects (e.g., vectors, balls, boxes, or cones) in continuous space and use geometric operations to simulate logical reasoning [2]. The intuition is that logical constructs such as subsumption, conjunction, and disjunction can be modeled using spatial relationships (e.g., containment or intersection). For example, EL Embeddings [39] represent concepts as Euclidean balls, where subsumption corresponds to containment. BoxE [35] uses axis-aligned boxes to capture more complex logical structures. Hyperbolic models such as ConeE [17] embed taxonomies as convex cones to preserve hierarchical structure. RotatE [57], though originally designed for knowledge graphs, uses complex rotations to capture properties like symmetry and inversion. These models offer strong logical expressiveness but are often computationally intensive and less suited for dynamic or user-defined query concepts.
In addition to vector-based methods, some models use algebraic structures such as lattices or order embeddings to capture subsumption hierarchies and logical entailment [62]. While expressive and theoretically grounded, these approaches are typically less practical.
The existing ontology embedding methods we reviewed above are designed for static ontologies with fixed concept sets and offer limited or no support for composing concepts using logical operations. For example, geometric models such as BoxE [35] and ConeE [17] can approximate conjunction using the intersection of regions. However, they do not support negation and offer limited or no support for disjunction. Since these methods require retraining to incorporate new concepts, they are not suited for dynamic, query-time composition. This makes them incompatible with our approach, where users define new composite concepts at runtime and embeddings must be computed on the fly without retraining. Our embedding method addresses this limitation by enabling on-the-fly construction of concept embeddings through fuzzy operations, supporting efficient and flexible concept-level search over large-scale ontologies.
2.2 Ontology Visualization and Exploration Tools
Ontology visualization tools are designed to help users explore, understand, and query complex ontologies. These tools typically fall into three categories:
Indented tree views display concept hierarchies in a collapsible format, making them intuitive for browsing taxonomies. Tools such as Kaon [27] support ontology editing and provide treemap-based navigation. WebProtégé [47], a widely used collaborative editor, also adopts a tree-based layout for concept exploration. OBO-Edit [19], designed for biomedical ontologies, uses a similar structure. While effective for viewing hierarchical relationships, these tools are less suitable for exploring complex inter-concept connections or non-hierarchical structures.
Graph-based visualizations, also known as node-link diagrams, represent ontology elements as nodes and their relationships as edges. This approach is better suited for capturing both hierarchical and non-hierarchical semantics. Tools such as OWLViz [45] and NavigOWL [34] extend Protégé with interactive graph views. TGVizTab [1], OWLPropViz [60], and OntoGraf allow users to dynamically expand and collapse nodes. WebVOWL [44] brings this functionality to the browser, providing interactive, user-friendly visualizations based on the VOWL specification. While graph-based tools reveal richer relational context, they often struggle with visual clutter as ontology size increases.
To improve scalability and clarity, several tools combine tree and graph layouts or adopt alternative visual metaphors. Jambalaya [42], a Protégé plugin, integrates hierarchical and relational views with enhanced layout control. PRONTOVISE [21] supports multiple visual modes, including bar charts, to compare concept properties. CODEX [32] focuses on ontology evolution and uses pie charts and word clouds to illustrate changes across versions. OntoTrix [8] replaces node-link diagrams with adjacency matrices to reduce clutter in dense graphs. More recently, Graffinity [55] has introduced a hybrid approach that combines matrix and node-link views to support the visualization of large ontologies.
In summary, tree-based tools are well-suited for browsing hierarchies, while graph-based and hybrid methods offer broader relational context. Although recent tools address scalability to some extent, visualizing and querying large ontologies in an interactive, user-friendly manner remains a challenge [22].
Most existing tools are primarily designed for navigating fixed hierarchies or inspecting predefined relations. They offer limited support for composing flexible, user-defined queries that involve multiple semantic conditions, such as conjunction, exclusion, or approximate similarity. When query functionality is available, it often relies on rigid form-based inputs or formal query languages that are difficult for non-expert users. Furthermore, many tools are built for lightweight ontologies with simple taxonomies and do not scale effectively to large, expressive ontologies with rich relational semantics. These limitations motivate the development of new tools that support intuitive and interactive construction of complex queries and enable semantic search across expressive, real-world ontologies.
3 Technical Background
This section provides the necessary background and preliminaries. We first review the basics of ontologies in Section 3.1, followed by an overview of fuzzy ontologies in Section 3.2.
3.1 Ontologies: Syntax and Semantics
In this work, we adopt the formalism of Description Logics (DLs) [7] to describe the syntax and semantics of ontologies, which is necessary for presenting our ontology embeddings. DLs are a family of logic-based languages that provide a formal, interpretable syntax and well-defined semantics for representing ontologies. They are widely used and form the foundation of popular ontology languages such as OWL [46], which combines DLs with the RDF data model [38]. Although our approach is not limited to DLs, we use DL notation to review fuzzy ontologies and their interpretations in order to clearly define our embeddings.
Formally, the syntax of a DL ontology consists of a signature , which defines the vocabulary of concept names, role (relationship) names, and individual names, and a knowledge base containing two types of axioms. The TBox () specifies relationships between concepts, including subsumption axioms (e.g., ), equivalence axioms (), and disjointness axioms (). The ABox () contains assertions about individuals, such as , indicating that individual is an instance of concept , or , stating that individuals and are related via the role . The semantics of an ontology is defined by an interpretation , where is a non-empty domain and is an interpretation function. This function maps each concept to a subset , each role to a binary relation , and each individual to an element .
Complex (or defined) concepts are constructed using logical constructors such as conjunction (), disjunction (), and negation (), which correspond to standard set operations in the domain. Specifically, the interpretation of a conjunction is given by , the interpretation of a disjunction is , and the interpretation of a negation is .
An ontology is said to be consistent if all of its axioms can be satisfied simultaneously in some interpretation. Consistency is a basic requirement for reasoning and querying, as inconsistencies can make the ontology logically meaningless or prevent useful inference. A concept is satisfiable if it has at least one instance in some interpretation; otherwise, it is considered logically contradictory. These notions provide the foundation for standard reasoning tasks such as subsumption checking, which determines whether holds in all interpretations (i.e., whether every instance of is also an instance of ), and instance checking, which verifies whether an individual necessarily belongs to a concept across all models of the ontology. While such reasoning tasks support classification and consistency checking, they are typically handled internally by DL reasoners.
In our setting, we are more concerned with querying ontologies to retrieve individuals or concepts that match a user’s information needs. Common query types include conjunctive queries (CQs), unions of conjunctive queries (UCQs), and recursive extensions that support reachability-like conditions. DL query languages such as SPARQL-DL [54], nRQL [30], and QueryTab [49] combine logical reasoning with structured query syntax. However, these languages require familiarity with formal syntax and detailed knowledge of the ontology’s structure, and their evaluation is often computationally expensive. This limits their suitability for users seeking intuitive, flexible, and interactive exploration.
3.2 Fuzzy Ontologies
Fuzzy ontologies extend classical ontologies by incorporating fuzzy logic, a many-valued logic in which truth values range continuously between 0 and 1. This enables modeling concepts and relationships that are inherently vague or uncertain. In contrast to crisp ontologies, where an individual either belongs to a concept or not, fuzzy ontologies allow individuals to belong to a concept with a certain degree of membership [10, 59]. For example, a patient might belong to the concept Diabetic with degree 0.8, reflecting partial evidence or uncertain data.
Fuzzy Description Logics (fuzzy DLs) extend classical DLs with fuzzy semantics, enabling formal reasoning over fuzzy sets [56, 31, 28]. As in the classical case, a fuzzy ontology has a signature and a knowledge base , but its interpretation function changes. Some fuzzy DL variants also introduce optional syntactic extensions, such as explicit degree annotations or fuzzy modifiers in axioms or assertions.
An interpretation now maps each concept to a membership function , where gives the degree to which individual belongs to . Similarly, roles are interpreted as fuzzy binary relations .
Complex concepts are interpreted using fuzzy operators. Given fuzzy membership functions and for concepts and , conjunction is defined as , disjunction as , and negation as , where is a continuous t-norm (e.g., product, Gödel, or Łukasiewicz), is its corresponding t-conorm, and is a negation function [25, 28]. These operators preserve the semantics of fuzzy set theory and allow the construction of defined concepts with graded membership.
Fuzzy DL also generalizes the notion of axiom satisfaction. A concept assertion is satisfied if the membership degree of in is at least , i.e., . A subsumption axiom is satisfied if, for all individuals in the domain, the degree to which implies is at least . This implication is computed using the residuum of the chosen t-norm, a standard fuzzy logic operation that captures graded entailment. An ontology is consistent to degree if all its axioms are satisfied to at least that degree.
In this work, we adopt a fuzzy Description Logic framework that supports partial membership in defined concepts and enables fuzzy reasoning over user-defined queries. Our implementation is based on fuzzy ontology reasoners and uses continuous t-norms to define fuzzy operations. This allows for flexible, approximate, and semantically grounded matching of complex concepts in expressive ontologies.
4 Ontology Querying using Fuzzy Ontology Embeddings
We now present our fuzzy ontology embedding approach and describe how it enables flexible, semantically meaningful queries over large ontologies.
Let be an ontology and let be a fuzzy interpretation of , where is a fixed set of domain elements and assigns to each concept a fuzzy membership function . The embedding of a primitive concept is defined as:
Embeddings for composite concepts are computed on demand using fuzzy logic operations applied element-wise to primitive embeddings. Let , , and denote the element-wise versions of a chosen t-norm, t-conorm, and negation operator, respectively. For example, the embedding of is , where could be, for instance, the product t-norm . Other common choices include the Gödel and Łukasiewicz t-norms. Disjunctions and negations are handled analogously using and , such as and .
A query is a composite concept that is defined compositional using primitive concepts and other composite concepts using fuzzy operators—conjunction (), disjunction (), and negation ()—without requiring these concepts to be explicitly present in the ontology. Each such query is represented by a vector , computed recursively from its structure using the fuzzy operators. Answering the query is done by retrieving similar primitive concepts. To retrieve relevant concepts, the system compares the embedding of a user-defined query with all primitive concept embeddings using cosine similarity and returns the top- most similar concepts. This enables exploratory, partial, and approximate matches that reflect semantic similarity rather than exact equivalence. Once a fuzzy interpretation is computed, the membership vectors for all primitive concepts can be precomputed and stored in a vector database. At query time, composite concept embeddings are computed dynamically using fuzzy operations, and similarity search is performed efficiently over the stored embeddings. This design enables fast, scalable, and semantically grounded concept-based retrieval without the need for repeated reasoning or re-embedding.
The quality of embeddings and the query results depend critically on the fuzzy interpretation used to assign membership degrees to domain elements. This interpretation must approximately satisfy the ontology’s axioms, ensuring semantic coherence. However, axiomatic satisfaction alone may not guarantee high-quality embeddings. For instance, embeddings should ideally reflect concept hierarchy, assign distinct vectors to semantically distinct concepts, and group semantically similar individuals with similar membership degrees. These additional criteria can be encouraged by integrating auxiliary axioms or constraints—e.g., enforcing similarity among sibling concepts or ensuring disjointness of non-overlapping concepts—even if such constraints are not part of the original ontology.
Our approach is agnostic to the specific fuzzy DL reasoner used to derive , as long as it yields consistent and interpretable membership functions. Several off-the-shelf reasoners can be used. FALCON [58] learns fuzzy interpretations via sampling and optimization. Users can specify the sample size, which defines the number of domain elements in the interpretation and thus determines the embedding dimension . This flexibility allows users to trade off between precision and computational cost. fuzzyDL [9] is a tableau-based reasoner that computes crisp or fuzzy satisfiability of axioms. It does not expose direct control over the size of the domain used for interpretation, making it less flexible for applications that require embedding generation over a fixed-size universe. DeLorean [48] approximates fuzzy reasoning using reduction techniques that map fuzzy ontologies into classical ones. It supports reasoning over various t-norms and fuzzy operators but similarly lacks user control over the interpretation domain. Among these, FALCON offers the most direct way to control the embedding space, as the number of sampled individuals directly determines the embedding dimensionality. This tunable design is well-suited for downstream applications that require balancing interpretability, query speed, and representational capacity.
Our querying approach and fuzzy ontology embeddings offer several advantages that make it well-suited for semantically grounded, scalable, and interactive ontology exploration:
Semantically grounded embeddings: The embeddings are derived from a fuzzy interpretation that satisfies the ontology’s axioms, ensuring that the resulting vectors reflect the intended semantics of the ontology. Unlike many prior embeddings that rely on taxonomic structures or shallow graph features, our method supports ontologies with expressive axioms, provided they are supported by the fuzzy DL reasoner. This allows meaningful semantic representation even in complex, non-hierarchical ontologies.
Compositionality and real-time querying: The embeddings are compositional: primitive concept vectors are precomputed once and stored offline, while composite concepts are constructed at runtime using efficient vectorized fuzzy operations. This enables real-time querying without retraining or re-embedding, in contrast to neural or graph-based approaches that require recomputation when new concepts are introduced.
Tunable embedding quality and size: The embedding dimensionality and fidelity are determined by the size of the interpretation domain , which can be controlled via parameters exposed by the fuzzy reasoner (e.g., the sample size in FALCON [58]). This provides a flexible trade-off between representational richness and computational efficiency.
Scalable and optimized retrieval: Since all concept representations are vectors, similarity search over primitive concepts can be delegated to optimized vector databases (e.g., FAISS [37], Qdrant, or Chroma). These systems support fast top- retrieval and efficient indexing, enabling responsive and semantics-aware retrieval for exploratory query interfaces.
4.1 Constructing Interpretations for Hierarchical Ontologies
In many practical domains such as biomedicine or botany, ontologies take the form of hierarchies or trees, where each concept has parent concepts, some number of children, and there are no cycles (e.g., HPO [29], PO [36]). For such ontologies, we can construct a synthetic but semantically meaningful fuzzy interpretation that captures hierarchical relationships and allows us to generate expressive embeddings efficiently.
We assume the ontology forms a tree-structured taxonomy, where all paths lead from leaves to the root through unique parent-child links. Let be the interpretation domain constructed iteratively as follows:
1.
At each step, a new domain element is added.
2.
A leaf concept is chosen uniformly at random.
3.
The membership of to is set to 1: .
4.
For every other leaf concept , we compute the graph distance , defined as the number of edges to their lowest common ancestor. We assign decaying membership using a fading parameter :
5.
For any internal (non-leaf) concept with children , the membership is defined recursively using the fuzzy union aggregation (e.g., probabilistic sum ):
This process generates a coherent fuzzy interpretation over a hierarchical ontology. By construction, the subsumption axioms are satisfied in the fuzzy sense, since the membership degree of any domain element to a parent concept is at least as large as that to its children, due to the properties of the fuzzy union. Thus, the generated satisfies all taxonomic axioms of the ontology.
This method provides an effective and interpretable way to construct embeddings from hierarchical ontologies, capturing both local concept similarity and the global structure. It also allows control over smoothness and generalization via the decay parameter , making it a practical tool for building fuzzy ontology embeddings in common tree-structured settings. We refer to these synthetic embeddings as -embeddings.
5 The FuzzyVis System
In this section, we present our prototype system, FuzzyVis. Section 5.1 describes its architecture and important implementation details. Section 5.2 further discusses the back-end server functionality. Section 5.3 does the same for the front-end interface. Lastly, section 5.4 explores the prior example 1 in greater depth as a usage scenarios that show how FuzzyVis supports ontology exploration and search through visual query building and approximate reasoning.
5.1 FuzzyVis Architecture and Implementation
FuzzyVis is a web-based system designed for ontology exploration. Figure 1 shows its overall architecture. The front-end is built using HTML5, CSS, and JavaScript, with D3.js (version 7) for interactive visualizations. The back-end is implemented in Python and communicates with the front-end via a FLASK API.
The system workflow begins with the user uploading ontology data in standard formats such as .owl or .ttl. The ontology is then processed and stored on the back-end. Once loaded, users can explore the ontology through interactive views and visualizations on the front-end. This exploration is supported by panels and controls that allow users to manipulate views, adjust parameters, and construct queries.
As users interact with the system, in the front-end FuzzyVis dynamically loads relevant ontology data and discards unused portions to optimize memory usage and ensure scalability for large ontologies. Many ontologies are so large that users cannot feasible view them in their entirety. By discarding unused portions users’ systems do not waste unnecessary resources. This optimization ensures that exploration of the ontology remains responsive. To maintain this responsiveness during querying, the back-end handles query resolution and returns the results to the query building panels of the front-end. Users can explore these results to navigate the ontology or refine subsequent queries.
While the front-end handles user interaction and visualization, all computations—including query evaluation and similarity search—are performed on the back-end. Embeddings for primitive concepts are stored in a Chroma vector database, 111https://github.com/chroma-core/chroma as described in Section 4. This enables quick query resolution, especially for ontologies with thousands of concepts, and or large embeddings.
5.2 Back-End Server
The back-end consists of four main components: the ontology loader, ontology-database, vector-database, query resolver. The ontology loader handles uploaded ontologies—pre-uploaded or submitted via the front-end. It parses ontology files and stores their contents in an ontology-database using the OwlReady2 python library [40]. During this process, additional metadata and statistics—such as subtree sizes and number of children—are computed and stored. The vector-database stores the embeddings for each primitive concept. The vector-database is separate from the ontology-database to perform quick similarity calculation and top-k retrieval. Our implementation uses Chroma for its vector-database. Just like for ontologies, users can submit embedding vectors or use pre-uploaded ones. Uploaded embedding vectors can be generated using whatever means users desire—such as fuzzy ontology reasoners—as long as the vectors have the properties of those we define in section 4. If users do not provide an embedding the system will generate a simple embedding using leaf concept membership and subsumption. The selection and loading of embeddings is also handled by the ontology loader. Lastly, the query resolver computes embeddings for user-defined composite concepts. The vector-database uses these composite concept embeddings to retrieve similar primitive concepts. A FLASK server API handles Communication between the back- and front-end. The front-end requests ontology information and the back-end provides.
5.3 Front-End Interface
The front-end is composed of several regions consisting of panels and controls that support interactive ontology exploration. Figure 2 shows an overview of the front-end interface as seen by the user. The primary visualization takes up the central region of the interface and is the users’ primary means of navigating the ontology (Figure 2.A). Users can change between different visualizations—treemps, nested lists, and network graphs—in this central region. The second region (Figure 2.B), the top header, provides quick access for concept search, switching between loaded ontologies, and loading ontologies. Users can manipulate the behavior and control the presentation using the visualization controls in the leftmost region (Figure 2.C). The first of the rightmost regions (Figure 2.D) is the concept panels where users can store concepts for future use and look at concept details. The second rightmost region (Figure 2.E) is the query building panels where users define queries and examine their results. FuzzyVis ’s front-end is compatible with all major web browsers.
Below, we describe the main components of the front-end interface in more detail.
5.3.1 Primary Visualizations
FuzzyVis ’s central region is contains the primary visualization (Figure 2.A). Here users see a representation of the ontology to aid them in exploration and understanding. Each component of FuzzyVis serves to modify, control, and navigate the primary visualization, with the back-end providing any information needed by the primary visualization. The primary visualization can take the form of any suitable representation of the ontology. Visualization should meet the following criteria:
Concept Distinguishability. Each concept in the visualization must be distinct and individually identifiable, at a reasonable scale. The density of concepts must be low enough that users can distinguish them. As well, concepts must have some clear identifier—often this is a label.
Consistent. The visualization should represent concepts and relationships in a consistent and identifiable manner.
Dynamic and Interactive. The visualization should react to users’ actions. Static representations may be sufficient for small and simple datasets, but ontologies are often neither small nor simple. Dynamic visualizations adapt to users needs and can change to show the ontology information that they are interested in—especially when presenting the entire ontology is impractical.
Property-revealing. The visualization should present secondary information and metadata of concepts. Features such as concept descriptions, depth, subtree size, and metadata contain information necessary to understand the content of ontologies. Users may be unfamiliar with concept labels and require secondary information to understand an ontology.
Space-efficient. The visualization should utilize the available space fully. Some ontologies have a large amount of content and visualizations that make poor use of space are either sparse or dense to the point of uselessness.
Structure-revealing. The visualization must showcase the primary semantic structure of the ontology. In most cases, this is the concept subsumption hierarchy. One of the primary purposes of ontologies is to express how concepts relate to one another. This information is critical to users’ understanding of ontologies.
Classical visualizations such as nested lists, icicle charts, network graphs, and treemaps can only partially meet these requirements. In fact, no single static visualization can satisfy every requirement. Network graphs often become unreadable with high volumes of concepts, nested lists poorly use space, and even treemaps struggle to convey secondary features. To address this problem, part of our solution is to supplement the primary visualization with secondary panels that provide information that the primary visualization cannot. The other part of the solution is to utilize dynamic interactive visualizations. To illustrate how we meet these requirements and explain the functionality of FuzzyVis we will discuss the nested treemap version of the primary visualization in this section.
Nested treemaps consist of nested rectangles, where each rectangle corresponds to a concept. Parent-child relationships are represented by the child concept’s rectangle being positioned inside the parent concept’s area. Nested treemaps are a suitable visualization of ontologies as they meet the majority of the requirements. In particular, such visualizations clearly distinguish concepts, are consistent, space-efficient, and structure-revealing. By extending nested treemaps to be dynamic and interactive, secondary properties can be shown and egregious concept density can be mitigated. An image of FuzzyVis ’s nested treemap, henceforth referred to as the treemap, is shown in Figure 3. As showing the entirety of ontologies is impractical, we implemented the treemap to only show concepts within an adjustable distance of a selected concept. As users interact with the treemap it adjusts to show the selected concepts, allowing users to navigate the ontology. Furthermore, the treemap has a number of additional features, interactions, and integration with the rest of FuzzyVis ’s front-end that aid users in their exploration activities.
Selecting concepts in the treemap refocuses the visualization onto them. By selecting concepts, users drill down into concept’s subtrees or up to through the parent—navigating the ontology. Beyond selection, users are able to drag concepts to other components, and pin concepts to mark them. Supplementary visual marks placed by users help them in future navigation by acting as signpost to aid in orientation. Furthermore, in cases where a concept is present in the subtrees of disjoint parents a visual mark present in multiple locations helps users identify this. For example, in HPO the Cardiac valve calcification concept is both a child in the Cardiovascular calcification and Abnormal heart morphology subtrees. When pinned Cardiac valve calcification shows up in both areas indicating to users overlap between the parent concepts. Without visual aids users need to manually read each concept label in order to identify them. Afterwards, once users have collected concepts that interest them, they can create queries using the Query Building Panels to direct their future exploration.
Users exploring ontologies may want to focus on a specific concept and see its content as though it were the ontology’s root. Simply increasing the depth quickly leads to two issues. One, child concepts a few layers deep have a small amount of space available to them; users must mouse over these children to see a tooltip to identify them. Two, when every subtree is full of child concepts the visualization becomes very dense, overwhelming users.
To address these issues, we implemented a toggle for a focus mode that enables on-the-fly visual distortions. During focus mode, when users mouse over concepts, FuzzyVis applies a discrete Cartesian distortion (often referred to as a fisheye distortion (or lens)), greatly increasing the moused-over concept’s area while maintaining its relative position to the rest of its siblings. Within the focused concept’s increased area, the subtree extends to show the children of the focused concept. Users that mouse over the added children can continuously drill down until there is insufficient space for further expansion. If users want to explore multiple subtrees simultaneously they can lock a concept as a locus and the distortion will maintain multiple focal points. Moreover, focus mode is not limited to extending the treemap deeper, as any focused locus can be transformed into a different visualization. For example, a concept cell can be converted to a node network graph or a panel showing class details.
As an example of the focus mode feature, in Figure 3, the depth is set to two and the Abnormal nervous system physiology is selected. Normally, no concepts other than the children of Abnormal nervous system physiology are visible under these conditions. Yet, with focus mode turned on, the user mousing over Abnormal central motor function fills the subtree with child concepts. Afterwards, when the user mouses over the child Abnormality of coordination further expands the subtree to show this concept’s children (e.g, Slurred speech). This could have continued until the user reached a leaf concept or ran out of screen space, but, the user decided this depth was sufficient and locked Abnormality of coordination as a locus. Thus, focus mode provides users with a dynamic and space-efficient way to quickly explore deeper into an ontology while maintaining context and their position. Furthermore, dynamic distortions are applicable to other kinds of visualizations, such as network graphs and charts.
Users control the presentation and behavior of the primary visualization through the Visualization Controls (discussed in detail in section 5.3.3). For the treemap, users can adjust the displayed depth, extend the labels with more information (e.g., depth numbers), control how concepts are tiled, which concepts are shown, and highlight sections of the visualization using these settings. Giving users control over the visualization options is necessary, as some ontologies present better under different conditions. For example, in HPO, among the children of Phenotypic abnormality there is Abnormality of the musculoskeletal system and Abnormality of the voice. Abnormality of the voice contains few classes and is small and almost unnoticeable in the treemap when cell scaling is set to be proportional to subtree size and no normalization is applied. Abnormality of the musculoskeletal system, on the other hand, takes up a significant amount of space. Users looking for conditions related to the voice may assume that HPO lacks them if they explored the ontology with the prior settings. With the cell scaling set to equal, each child concept of Phenotypic abnormality takes up the same amount of space. Users see both Abnormality of the musculoskeletal system and Abnormality of the voice and are less likely to miss important concepts. Furthermore, users can toggle these options to see how the visualization changes—toggling the equal scaling on and off shows users which concepts take up the majority of an ontology while quickly returning to a view suitable for navigation. The ability to adjust the visualization lets users see more of an ontology’s information without restricting them to suboptimal visualizations.
FuzzyVis’s primary visualization helps users to explore the structure and semantics of ontologies. In conjunction with other front-end panels, it provides a high-level spatial overview that enables users to identify key concepts and regions of interest. By presenting concepts with a dynamic visualization FuzzyVis supports rapid navigation aiding in users mental model formation. This visual aid is necessary when working with complex knowledge domains where manual inspection of concepts would be inefficient or overwhelming. Ultimately, the primary visualization serves as the main interface through which users interact and explore the ontology.
5.3.2 Header Bar
FuzzyVis’s header bar (Figure 2.B) has three interactive elements. Users can load ontologies, switch between loaded ontologies, and perform keyword search to quickly navigate using these interactive elements. From left to right, it includes a search bar for concept lookup, tabs for navigating between ontology instances, and a plus button to open the ontology selection panel (Figure 4).
The first header element is the search bar. Users can look for concepts by their labels and directly jump to them. Users’ input text is matched against concepts labels in the current system. Expert users already familiar with an ontology can quickly navigate to known concepts using the search bar. As well, users can check whether a concept exists in a loaded ontology. For instance, in a biomedical ontology, users can search for Heart to verify its presence. If no results are returned, the concept is likely absent. Related results, such as Heart block, are also shown, helping users discover similar concepts. The search bar is not FuzzyVis’s primary means of navigating the ontology and is instead intended for quick repositioning to known landmark concepts.
The second header element consists of tabs to switch between multiple loaded ontologies. Each loaded ontology retains its own settings, highlight rules, pinned concepts, queries, and visualizations. Users can explore multiple related ontologies in parallel, or explore the same ontology using different embeddings and fuzzy operators. The tabs allow users to organize their queries into separate instances, create dedicated highlighting rules for different ontology regions, and compare the effectiveness of different embeddings. For large ontologies, having multiple instances of an ontology helps users explore different regions without having to remember exactly where they were or repeatedly update the concept collection. Users add new tabs with the ontology selection menu.
The last header element opens the ontology selection menu (Figure 4. Users can choose from preprocessed ontologies or upload their own. Uploaded files are sent to the back-end for processing. The back-end stores each ontology and its embedding in separate temporary databases. Preprocessed ontologies each reuse their own static database. The embedding options act the same as the ontology ones. Users can upload different embeddings to evaluate their effectiveness or use the pre-made ones. The last options are for configuring the fuzzy logic operators used for query resolution. Depending on the fuzzy DL reasoner used aligning an embedding with the matching fuzzy operators may be necessary. The options of product, Gödel, or Łukasiewicz t-norms cover the majority fuzzy reasoners and even when mismatched with the reasoner either of the three will often work.
5.3.3 Visualization Controls
The left region of FuzzyVis contains controls and panels for manipulating the primary visualization (Figure 2.C). To illustrate, panels for the treemap visualization are shown in Figure 5 This region holds three components. First, the highlight panel, for setting rules to stain the primary visualization with colors (Figure 5(a)). Second, the visualization control panel, for manipulating the behavior of the primary visualization (Figure 5(b)). Last, is a help panel to explain to unfamiliar users how to interact with the system. The left region’s purpose is to allow users to tune the visualization to more closely meet their needs. As users need vary, a single set of static visualization parameters is often insufficient. For example, a user with a smaller monitor may require a more horizontal tiling for a treemap in order to be able to read the labels without mousing over the cells.
The highlight panel (Figure 5(a)) allows users to annotate the primary visualization. In this panel, users construct rules and assign colors to highlight ontology concepts. Concepts which if they satisfy multiple rules will mix the highlight colors. Users can filters based on concept label, subtree size, number of children, or other criteria when making rules. Rules that restrict on string values can either require an exact match or just substring matching. Rules that restrict on numerical values can set minimums, maximums, equality, and ranges. Additionally, users can negate rules or toggle them on and off.
For example, a highlighting rule may specify: color blue all concepts with labels that contain bulbar (the first rule in Figure 5(a)). This rule stains concepts such as Bulbar Palsy and Pseudobulbar signs blue making them significantly more noticeable. Highlighting allows users to quickly identify concepts of interest to form landmark and neighborhood knowledge. Isolated colored concepts exhibit uniqueness, while groups of similarly colored concepts could bear similarity to each other depending on the rules. Furthermore, users can set in the controls to only display highlighted concepts in the primary visualization. For large ontologies with many concepts some visualizations will densely pack concepts making it hard to distinguish them (ex. treemaps, or network graphs). In such scenarios, highlighting provides needed visual marks for distinguishing concepts. Moreover, this panel’s function can be expanded to support ontology-specific rules, depending on an ontology’s properties.
The visualization control panel is used to configure the behavior of the primary visualization. Here users can adjust the parameters that determine how the primary visualization is displayed. Depending on the selected visualization options will differ. For example, Figure 5(b) is the controls for a nested treemap visualization. For such treemaps, settings include:
Depth. The max displayed depth from the selected concept. Higher values will densely pack the interior of cells and allow quick traversal to lower layers. Where as, lower values improve clarity.
Tiling Ratio. The shape of the treemap cells. User preference for more or less horizontal and vertical cells is controlled with this setting. Depending on screen size and ontology content some ratios can improve or reduce clarity significantly. For example, a horizontal ratio may make long labels easier to read, but with too many children concepts there many not be enough vertical space to show the label when all concepts are stacked as plates.
Scaling. The value that determines the size of concepts. Users interested in metadata such as subtree size or number of children can choose to scale concepts based on these values. A very large concept contains the majority of the ontology and can be considered more important. On the other hand, equalizing scaling lets users see concepts that would be hidden due to the dominance of others.
Visibility. Which concepts are displayed. Based on highlighting rules determine which concepts are shown in the primary visualization. Users only interested in specific concepts can hide those they deem irrelevant.
Normalization. Adjust the scaling so that minor concepts are more visible and dominate ones take up less space. Another means by which users can control how concepts are displayed.
Focus Mode. Enable mouse over distortion of the visualization. Users can dynamically adjust the scaling of concepts to focus on what interests them regardless of the ontology’s structure.
Showing Layer Numbers. Annotate concepts with their layer number. Some users may struggle to distinguish depth in some visualizations. For example, in a network graph it can be difficult to distinguish how deep a concept is in a given subtree.
Toggling similarity stain. Stain the concepts with colors based on the query results. Users exploring queries can quickly see what parts of the ontology correspond to their query.
Some of these settings are universal, others, are visualization specific. In general, the purpose of these controls is to align the visualization to users’ needs. As ontologies can have a great amount of variety, default settings are not guaranteed to be optimal—allowing users to adjust the settings obviates this issue.
The last panel, help and keybindings provides users with information on how the system functions. Users can hide this view if it is unneeded. Though a full interactive tutorial would be more optimal it is beyond the scope of this prototype.
5.3.4 Concept Panels
The first of the right region of FuzzyVis are the concept panels (Figure 2.D). These panels have two primary functions: organize a collection of concepts for future use and provide concept details. As users explore ontologies, they encounter concepts of interest that they may want to revisit or use in queries. These concepts may be located in disparate regions making them inconvenient to access. As well, users need concept details to evaluate relevance and suitability. In HPO, users investigating skull-related conditions may be interested in the Abnormal cranial nerve physiology and Abnormality of the head concepts. Yet, neither of these concepts are conveniently close to each other. Nor without seeing details can users confirm that these concepts truly meet their requirements. When users store these concepts in the collection they can quickly navigate to and between them. Overall, the concept panels are present to aid users in their navigation and query formulation.
The collection panel (Figure 6(a)) stores concepts of interest for easy access. Users drag and drop, use any of the collection keybindings, or pin concepts to add them to the panel. Additionally, users control which concepts are pinned here. The primary visualization annotates with a visual mark any pinned concepts for easier identification. The concepts within the collection can be dragged to any other component that takes concepts as inputs. The highlight component (Figure 5(a)) constructs a rule to match dropped concepts. The primary visualization refocuses onto the dropped concept. As well, when dropped on, any query being built will have the concept added to it. Users’ collections allow them to quickly navigate back to concepts of interest and not have to remember the full names of every concept that they are interested in. A user selecting Abnormality of the head can relocate the primary visualization there regardless of where they currently are exploring.
The details panel (Figure 6(b)) shows information about the currently selected concept. Dependent on the ontology there may be more or fewer entries. However, common among most ontologies are values for: label, IRI, parent, and children (concepts which possess a subclass relationship with a given concept). Thus, this panel can show a significant amount of critical information. For example, HPO includes concept definitions which provide clarity as to what a concept represents. While the primary visualization provides high-level spatial information for the ontology concepts, it forgoes details in favor of visual clarity. As such, the details panel is necessary to provide low-level detailed information to users for individual concepts.
5.3.5 Query Building Panels
The second section of FuzzyVis ’s right region are the query building panels (Figure 2.E). Users interact with these components to create, resolve, and analyze queries. Hierarchical exploration of ontologies exposes users to only a portion of the semantics and structure. Furthermore, analyzing individual concepts one at a time is inefficient. Thus, as users explore they may want to verify their inferences or locate unknown, yet identifiable, concepts. The query building panels enable users to perform such non-linear exploration and develop an understanding of the structure and semantics of ontologies. We discuss this in-depth in our usage scenario (Section 5.4).
Users construct their queries with the query builder control (Figure 7(a)). Queries are composed of concepts, other queries, and support standard logical operators (AND, OR, and NOT). These operators correspond to intersection, union, and negation respectfully. When users finalize their queries, FuzzyVis ’s front-end passes them to the back-end query resolver which converts them to composite concepts. These composite concepts have their embeddings compared to the ontology’s primitive concepts before a list of top-k most similar concepts are returns. There is no restriction on which concepts can be added to a query, nor is there anything stopping a user from creating a logically inconsistent statement. Such a query is useless when resolved with traditional embeddings. However, our fuzzy ontology embeddings can still capture semantic information from some such queries.
The query results are shown in FuzzyVis ’s query results panel (Figure 7(b)). Here users can see the degree to which each primitive concept is similar to their query’s composite concept. Users can expand any of the results to show concept details. As well, users can drag the concepts from the results list and use them in future queries or in other regions of the system. The similarity stain option in the visualization control panel works with these results. When this option is toggled on the similarity values of the query results stain the primary visualization with a matching tone. The query resolver uses cosine similarity to determine the similarity score. Because FuzzyVis stores primitive concept embeddings in a vector database queries resolve nearly instantly.
5.4 Usage Scenario
The following usage scenario demonstrates the features of FuzzyVis and the practical application of fuzzy logic queries for ontology exploration. We use the HPO for this scenario. We assume the user knows basic medical terms but is unfamiliar with HPO. The embedding used is an -embedding (as described in Section 4.1) with . For fuzzy operators, we use the product t-norm for conjunction, its t-conorm for disjunction, and the standard negator for negation.
A physician, henceforth referred to as the user, has a patient who has difficulty speaking and swallowing. This patient has no history of immune disorders and tests reveal no signs of infection. To help the patient, the user wants to accomplish two goals. One, they want to identify possible conditions that could explain the patient’s symptoms. Two, they wish to learn about any secondary ailments that the patient may experience now or in the future. The user looks to HPO as it contains information about abnormal conditions. However, the user needs a tool to assist with exploring the ontology due to its complexity and size. Thus, the user launches FuzzyVis to explore HPO in order to help their patient.
The user opens FuzzyVis ’s front end in their browser of choice and loads HPO with the previously mentioned parameters. After initializing, FuzzyVis is ready to use and defaults the primary visualization to a nested treemap. The user sees the initial state of the primary visualization and adjusts the settings in the visualization control panel. First, the user sets the scaling of each cell to be equal as the relative size of concepts is not relevant to their goals. Second, the user sets the tiling so that the concepts clearly show their labels. Lastly, they set the depth to a low value as HPO has many terms and showing multiple layers is overwhelming. Now each concept is clearly visible to the user. Next, the user creates a few highlight rules for key terms they are interested in. For this example, the highlight rules are to stain any concept with the words voice, speech, and swallowing in their label for easier identification. After this set up the user can now quickly identify initial concepts of interest. Users of FuzzyVis can with little effort setup the primary visualization in ways that streamline their future searches.
Seeing the concept Phenotypic abnormality, the user selects it and begins their search. During their search, the user identifies several child concepts related to their patient’s condition. These concepts are pinned (for easier identification) and added to the concept collection for use in queries. To explore more specific concepts, the user activates focus mode and starts browsing the contents of Phenotypic abnormality’s children. They see that Abnormality of the voice does not contain what they are looking for; the user collects it for query use regardless. Additionally, the user spots Abnormality of the immune system and adds it to the collection so it can be negated in future queries. Hoping to quickly find something, the user interacts with the search bar and looks for concepts whose labels contain the word speech. Among the keyword search results is the concept Slurred speech, a direct match for the user. Selecting Slurred speech takes the user deeper into the ontology, where they then pin it. Wanting to know where in the ontology they are, the user toggles on the depth numbers and sees that Slurred speech is eight layers deep and has no children. Then, the user increases the depth to view the full ancestry and spots that Slurred speech is within the Abnormality of the nervous system subtree. Now that the user knows where to look, they select Abnormality of the nervous system, reduce the shown depth. and continue their search here. Already the user has started to identify landmark concepts related to their problem, has learned the routes between them, and is now aware of the neighborhood in the ontology that contains concepts of interest. FuzzyVis allows for fast and seamless exploration; users can quickly see concept contents at a glance and freely navigate ontologies.
The user’s exploration continues for only a bit longer. Settling on Abnormality nervous system physiology the user encounters the concept Dysphagia within it. The focus mode alternative views and the concept details panel allow the user to quickly see the definition of this concept. As Dysphagia is the medical term for difficulty in swallowing, and with the prior discovery of Slurred speech, the user is now confident that they can start constructing queries. The user formulates their query as the presence of Slurred speech and Dysphagia while excluding Abnormality of the immune system. Using the query building panels the user drags concepts from the collection to the panel to create their query. Ultimately, the user defines the composite concept :
The user sends in their query, FuzzyVis ’s back-end quickly processes it, and the front-end updates nearly instantly with results. Among these results are concepts such as Pseudobulbar paralysis, Pseudobulbar signs, and Abnormal esophagus physiology. Wanting to locate these concepts in the ontology, the user turns on the similarity stain and spots the Pseudobulbar signs concept. With focus mode on, the user locks and splits Pseudobulbar signs to show its details and expands it to show Pseudobulbar paralysis. The results of both the similarity stain and this focusing can be seen in Figure 8. The pseudobulbar concepts refer to neurological conditions that explain the patient’s symptoms. On the other hand, Abnormal esophagus physiology is an alternative direction for investigation as it can also explain the symptoms. These results have now narrowed down potential causes that the user can investigate to diagnose their patient. Finding these concepts with keyword search would have been difficult for the user as their vocabulary did not perfectly match the ontology’s vocabulary. Moreover, the user would not know how to formulate a full SPARQL query to find these concepts. Users of FuzzyVis do not need to write out queries, instead, they build them through interactions with the interface by dragging and dropping concepts and other queries to define complex concepts. FuzzyVis ’s visual query building obviates the need for users to precisely spell out concept labels and by supporting nesting queries helps users separate the components of their queries into understandable segments. In allowing users to create fuzzy queries, FuzzyVis lets users find concepts relevant to their needs even when their requests are imprecise.
An example of the imprecise kinds of queries possible is: our scenario user’s interest in finding secondary ailments, a task they did by updating their query to exclude Abnormality of the voice. This would be an illogical query. However, because the -embedding contains fuzzy membership across the ontology the user believes that this query may still offer useful insights. Thus, the user sends in the updated query :
returns concepts related to digestive issues. This is expected as Dysphagia is part of the Abnormal esophagus physiology subtree which belongs to Abnormality of the digestive system. Thus, the condition of the patient may have consequences for the digestive system. The fuzzy ontology embeddings we use capture such relationships and continue to preserve them in complex composite concepts created by users. Armed with this new knowledge the user can ask their patient further questions to narrow down their condition. To check for Pseudobulbar signs they can inquire if the patient has found themselves more irritable and prone to emotional outbursts. To check for Abnormal esophagus physiology the user can simply inspect the patient’s throat.
This usage scenario highlights how FuzzyVis ’s features and fuzzy ontology embeddings can be used to explore ontology content. Despite having no prior knowledge of HPO’s structure and content, or query languages like SPARQL, the user was able to find relevant concepts and develop an understanding of the ontology. The system’s interface—composed of interactive visualizations—empowers users to visually build queries and explore the ontology. Furthermore, the use of fuzzy logic enables users to create queries even without expert knowledge of the ontology. These features of FuzzyVis and fuzzy ontology embeddings represent a way of exploring ontological information that is more flexible and user-friendly.
6 Conclusion and Future Work
This paper introduced FuzzyVis, a prototype system for visual query building over ontologies using fuzzy logic and membership-based embeddings. FuzzyVis allows users to define complex, possibly vague concepts using familiar set operations such as intersection, union, and negation. These user-defined concepts are interpreted as fuzzy sets and embedded in a continuous vector space, enabling efficient similarity-based query evaluation. The system supports exploratory ontology search through an interactive visual interface that combines search, navigation, query construction, and result presentation. FuzzyVis demonstrates how fuzzy embeddings can enhance user interaction with large ontologies, making query formulation more accessible and intuitive, especially for non-expert users. Unlike traditional querying methods that rely on formal syntax or strict logical reasoning, FuzzyVis bridges symbolic and sub-symbolic techniques to support approximate reasoning and flexible concept retrieval.
FuzzyVis is a prototype system, and there are several directions for future development. First, improving system performance and scalability for very large ontologies remains a priority. This includes optimizing both front-end visualizations and back-end embedding computations. Second, expanding FuzzyVis’s features can enhance its usability. For example, incorporating additional visualizations such as collapsible tree views can support different user needs. Similarly, adding new views for triaging and organizing query results would improve the analysis and comparison of retrieved concepts. Third, exploring alternative embedding techniques—especially those that are more scalable or tailored to specific ontology types—may lead to better performance or interpretability. Finally, an important research direction is to extend FuzzyVis to support more expressive ontologies with richer axioms beyond hierarchical subsumption structures. This would require augmenting both the query builder and embedding mechanism to handle a broader range of logical constructs and relationships.